A beaker containing 200g of water. The heat capacity of the beaker is equal to that of 20g water. The initial temperature of water in the beaker is 20°C. If 440g of hot water at 92°C is poured in,the final temperature(neglecting radiation loss) will be nearly

Grade:12

1 Answers

MPRANEETH
7 years ago
we knowthat principle of calorimetry :heat lossed by the hotbody is equal to heat gained by the cold body
there we know that Q=mc(dt)
therefore equating to theheat gained by cold body
given total mass is (200+20)(T-20)=440*((92-T))
THEREFORE  ON SIMPLIFYING WE GET T =68
iassumed the final temperature T